if sum of age of three brothers are 42 and before 10 years ratio of their ages are 1:2:3 then what is the age of younger brother.?
Answer by ankor@dixie-net.com(22740)

before 10 years ratio of their ages are 1:2:3
let x = the multiplier ten yrs ago
then 10 yrs ago, their ages:
1x, 2x, 3x
Now
if sum of age of three brothers are 42
(x+10) + (2x+10) + (3x+ 10) = 42
6x + 30 = 42
6x = 42 - 30
6x = 12
x = 12/6
x = 2 is the multiplier
:
then what is the age of younger brother.?
2 + 10 = 12 yrs is the age of the younger brother
:
:
see if the adds up, 10 to their ages 2, 4, 6
12 + 14 + 16 = 42